Grade 12Physical Chemistry

eaction of first order after 100 min 75g of reactant a was decompose when 100g is initial taken calculate time required when 150g of a was decompose

Aniket
9 Years agoGrade 12

2 Answers

Vikas TU
9 Years ago
From the first order reaction system we know that,
K = [2.303log(A0/A)]/t
k = [2.303log(100/25)]/100
For 150 g,
[2.303log(100/25)]/100 = [2.303log150/75]/t
0.6/100 = 0.3/t
t = 0.02 s
 
Anjali
6 Years ago
Chin cchino fixups costars opcodes Cocteau l claro pass Claiborne Oxon Ixia ALC concisely Paulson claw will come Pomona's servicing mixtape Khush knows CCOl luck KVP logic Crawley mom marram
Vivian
Gibbons AQA apricot